OpenStudy (anonymous):

Solve for x x2 - 3|x - 2| - 4x = - 6

myininaya (myininaya):

we have two cases to consider

myininaya (myininaya):

since |x-2|=x-2 when x-2>=0 and |x-2|=-(x-2) when x-2<0

myininaya (myininaya):

you have two equations to consider and solve \[x^2-3(x-2)-4x=-6 \text{ or } x^2+3(x-2)-4x=-6\]

myininaya (myininaya):

both of these are just quadratic equations that need fixing up

OpenStudy (anonymous):

Thankyou.